A penny is tossed and a 6-sided fair die is rolled. What is the probability of tossing a tail and rolling and even number?
A. 1/2
B. 1/12
C. 1/4
D. 5/12

Answer by stanbon(57332)

A penny is tossed and a 6-sided fair die is rolled. What is the probability of tossing a tail and rolling and even number?
These are independent events.
P(tail) = 1/2
P(even) = 1/2
---
P(tail and even) = 1/4
=======================
Cheers,
Stan H.
=======================
A. 1/2
B. 1/12
C. 1/4
D. 5/12